Two glass vessels M and N are connected by a closed valve.
MN
M contains helium at 20 °C at a pressure of 1 × 105 Pa. N has been evacuated, and has three times the volume of M. In an experiment, the valve is opened and the temperature of the whole apparatus is raised to 100 °C.
What is the final pressure in the system?
1
Expert's answer
2016-05-26T13:28:03-0400
pV/T = p_1V_1/T_1
V_1 = 4V
p1 = pVT_1/V_1T = PVT_1/4VT = pT_1/4T = 1*10^5 * (373.15+20) / 4* (373.15+100) = 2.07*10^4 (Pa)
Answer
p1 = 2.07*10^4 (Pa)